### v1

@@ Abstract
Summary:— We calculate the $I^{G}J^{PC}=0^{+}1^{++}$ channel $D\bar{D}^{*}$ scattering at four pion masses ranging from 250 MeV to 417 MeV in $N_{f}=2$ lattice QCD. We use the distillation method to calculate the related correlation matrix of $c\bar{c}$ operators and $D\bar{D}^{*}$ operators. At $m_{\pi}\approx 417$ MeV where there are no severe OPE lhc issues, a shallow bound state is observed to reside below the $D\bar{D}^{*}$ by $1.3_{-1.0}^{+0.8}$ MeV, and is a $D\bar{D}^{*}$ component of 96(1)%. When considering the OPE lhc effect, the results at the lighter pion masses also indicate the possible existence of a bound state. In the meantime, the finite volume energy levels above the $D\bar{D}^{*}$ threshold hint the existence of a $1^{++}$ resonance below $4.0$ GeV with a width around $60$ MeV. This is compatible with the expectation of phenomenological studies.
@@
Acknowledgements:— We are grateful to Profs. Qiang Zhao, Feng-Kun Guo, Jia-Jun Wu and Xu Feng for the valuable discussions. This work is supported by the National Natural Science Foundation of China (NNSFC) under Grants No. 11935017, No. 12293060, No. 12293061, No. 12293065, No. 12293063, No. 12075253, No. 12070131001 (CRC 110 by DFG and NNSFC), No. 12175063, No. 12205311, the National Key Research and Development Program of China (No. 2020YFA0406400) and the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030302). The Chroma software system Edwards and Joo (2005) and QUDA library Clark et al. (2010); Babich et al. (2011) are acknowledged. The computations were performed on the HPC clusters at the Institute of High Energy Physics (Beijing) and China Spallation Neutron Source (Dongguan), and the ORISE computing environment.
S1. Conventions and Operators
@@
$\begin{cases}|u\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |d\rangle=\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}\begin{cases}|\bar{d}\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |\bar{u}\rangle=-\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}$ (SI1)
@@
$\begin{cases}|D^{(*)+}\rangle=|c\bar{d}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)0}\rangle=|c\bar{u}\rangle=-|\frac{1}{2},-\frac{1}{2}\rangle,\\ |\bar{D}^{(*)0}\rangle=|u\bar{c}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)-}\rangle=|d\bar{c}\rangle=|\frac{1}{2},-\frac{1}{2}\rangle.\\ \end{cases}$ (SI2)
@@
$|D\bar{D}^{*}\rangle_{I=0}^{Q=0}=\frac{1}{2}\left(|D^{+}\bar{D}^{*-}\rangle+|D^{0}\bar{D}^{*0}\rangle-|\bar{D}^{0}D^{*0}\rangle-|D^{-}D^{*+}\rangle\right).$ (SI3)
Since the charge conjugation transformation ($\mathcal{C})$ of $D$ and $D^{*}$ are conventionally defined as $\mathcal{C}|D\rangle=+|\bar{D}\rangle$, $\mathcal{C}|D^{*}\rangle=-|\bar{D}^{*}\rangle$, one can easily check the $\mathcal{C}$ transformation $\mathcal{C}|D\bar{D}^{*}\rangle_{I=0}^{Q=0}=+|D\bar{D}^{*}\rangle_{I=0}^{Q=0}$. The $\mathcal{P}$-parity $P=+$ requires the $D\bar{D}^{*}$ relative angular momentum $l=0$. Therefore, the $S$-wave $D\bar{D}^{*}$ state with the combination in Eq. (SI3) has the desired quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. The combination in Eq. (SI3) gives also the flavor structure of the interpolation field operator of $D\bar{D}^{*}$ systems.
A $S$-wave two-particle operator $\mathcal{O}_{AB}(q)$ is easily constructed by the summing over all the relative momenta $\mathbf{p}\equiv 2\pi\mathbf{q}/L$ with the same $q=|\mathbf{q}|$, namely,
$\mathcal{O}_{AB}^{q}=\frac{1}{N_{q}}\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q}).$ (SI4)
where $R$ runs over all the elements of the octahedral group $O$ and $N_{q}$ is the degeneracy of $q$. Given the single-particle operators of $D$ and $\bar{D}^{*}$, $(\mathcal{O}_{D},\mathcal{O}_{D^{*}})\equiv(\bar{u}(\bar{d})\Gamma_{D}c,\bar{u}(\bar{d})\Gamma_{D^{*}}c)$, we combine the relations in Eq. (SI3) and (SI4) to build up the $D\bar{D}^{*}$ operators $\mathcal{O}_{D\bar{D}^{*}}^{q}$ with the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. In practice, we use two combinations of $\Gamma_{D}$ and $\Gamma_{D^{*}}$, namely, $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$ and $(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$. The operators $\mathcal{O}_{D\bar{D}^{*}}^{q=0,\gamma_{5}(\gamma_{4}\gamma_{5})}$ with the two $(\Gamma_{D},\Gamma_{D^{*}})$ combinations (indicated by the superscripts $\gamma_{5}$ and $\gamma_{4}\gamma_{5}$) are also named by $\mathcal{O}_{4}$ and $\mathcal{O}_{5}$, respectively. The $\mathcal{O}_{D\bar{D}^{*}}^{q=1}$ operator uses the first $(\Gamma_{D},\Gamma_{D^{*}})$ combination and is also named as $\mathcal{O}_{6}$.
@@
$\mathcal{O}_{c\bar{c}}^{r}(t)=\frac{1}{N_{r}}\sum\limits_{|\mathbf{y-x}|=r}\bar{c}(\mathbf{x},t)\gamma_{5}\gamma_{i}K_{U}(\mathbf{x},\mathbf{y};t)c(\mathbf{y},t),$ (SI5)
@@
Figure SI1: Schematic quark contraction diagrams for the calculation of the correlation matrix $C_{\alpha\beta}(t)$. For convenience, the operator set is sorted as $\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\}$. All the contributions are considered except for charm quark annihilation diagrams which are expected to be suppressed by the OZI rule.
@@
Figure SI2: Effective mass using different operators subset and solving GEVP, ranging from different $m_{\pi}$. In each column of the panels, the headers indicate the operators involved in GEVP. The colored lines show the effective mass behavior corresponding to their dominant operators, as indicated by the same color. Their fitting values are given in Table.[SI1-SI4].
S2. Determination of finite volume energies
Based on the operator set introduced above,
$\mathcal{S}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\},$ (SI7)
we calculate the corresponding correlation matrix
$C_{\alpha\beta}(t)=\frac{1}{T}\sum\limits_{\tau}\langle\mathcal{O}_{\alpha}(t+\tau)\mathcal{O}_{\beta}(\tau)\rangle~{}~{}(\alpha,\beta=1,2,\ldots,7),$ (SI8)
where the correlation functions with different source time slices $\tau$ are averaged to increase the statistics. Figure SI1 shows the quark diagrams after Wick’s contraction that are involved in the calculation of $C_{\alpha\beta}(t)$. All the connected quark diagrams and the diagrams including the light quark annihilation effects are taken into account, while the charm quark annihilation effects are neglected owing to the OZI rule.
    $C_{ij}(t_{1})v_{j}^{(n)}(t_{1},t_{0})=\lambda^{(n)}(t_{1},t_{0})C_{ij}(t_{0})v_{j}^{(n)}(t_{1},t_{0}),$ (SI9)
    for given $t_{1},t_{0}$. Thus we obtain the optimized correlation functions $C^{(n)}(t)=v_{i}^{(n)}(t_{1},t_{0})v_{j}^{(n)}(t_{1},t_{0})C_{ij}(t)$. This procedure runs over the cases at the four $m_{\pi}$’s. The effective masses of $C^{(1)}(t)$ and $C^{(2)}(t)$ are illustrated in the rightmost column (column (e)) in each panel of Fig. SI2. When $\mathcal{O}_{7}(J/\psi\omega,q=0)$ is added to the operator subset, the solution to the corresponding GEVP gives an additional energy level close to the $J/\psi\omega$ threshold, as shown in column (c) as black points in each panel of Fig. SI2.
  • Now we perform the GEVP analysis to the operator subset $\{\mathcal{O}_{4},\mathcal{O}_{5},\mathcal{O}_{6},\mathcal{O}_{7}\}$. The effective masses of the four optimized correlation functions are plotted in column (d) of each panel of Fig. SI2, where the effective mass illustrated by black points is almost the same as that in column (c). Other three effective masses have good signal-to-noise ratios but do not show plateaus even in the time range $t/a_{t}\in[25,40]$.
  • If the GEVP analysis is performed to the full operator set $\mathcal{S}$, we obtain seven optimized correlation functions. The effective masses of the lowest six states are plotted in column (a) of each panel of Fig. SI2. One can see that the effective masses of the lowest five states tend to saturate on plateaus. Column (b) in each panel of Fig. SI2 shows the effective masses when the $J/\psi\omega$ operator is excluded from the operator set. In comparison with column (a), the state illustrated by black points disappears but other states are still there and almost do not change.
The observations above indicate that the $J/\psi\omega$ operator ($\mathcal{O}_{7}$) almost does not couple with $c\bar{c}$ and $D\bar{D}^{*}$ operators and are therefore nearly irrelevant to the $D\bar{D}^{*}$ scattering. In contrast, $c\bar{c}$ and $D\bar{D}^{*}$ operators strongly couples with each other and must be considered together. These observations are in agreement with the previous work Ref. Prelovsek and Leskovec (2013).
As for the energy levels $E_{n}$ to be determined, we take the convention that the energy levels reflected by the effective masses in column (b) of Fig. SI2 are ordered as $E_{1},E_{2},\ldots,E_{6}$ from bottom to top. For a specific subset of the full operator set $\mathcal{S}$, the energy level that is adjacent to $E_{n}$ has the same index $n$. We do not include the energy level (indicated by black points) corresponding to the $J/\psi\omega$ to the level ordering.
For each specific operator subset, we first perform two-exponential fits to the optimized correlation functions $C^{(n)}(t)$ introduced above and take the lowest energy to be an estimate of $E_{n}$, which are listed in Tables SI1-SI4 for all the four $m_{\pi}$’s. In the data blocks in each table, the first two rows of the leftmost block are almost the same and are the energy levels $E_{n}$ from $c\bar{c}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in first two rows of the middle block are from $D\bar{D}^{*}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in the last two rows are from both $c\bar{c}$ and $D\bar{D}^{*}$ operators (with and without the $J/\psi\omega$ operator. The energy values in bold are determined from the ratio method (see below). It is clearly seen that the inclusion or exclusion of the $J/\psi\omega$ operator almost does not affect other energy values. So we exclude the $J/\psi\omega$ operator in the following discussions.
Comparing the energy levels from only $c\bar{c}$ operators, those from only $*D\bar{D}*$ operators, and those from both $c\bar{c}$ and $D\bar{D}^{*}$ operators, the strong correlation between $c\bar{c}$ and $D\bar{D}^{*}$ operators is obviously manifested. Especially, the three energy levels from only $*D\bar{D}*$ operators are far from non-interacting $D\bar{D}^{*}$ energies. This is attributed to the light quark annihilation effects illustrated in Fig. SI1. Therefore, it is necessary to include $c\bar{c}$ operators when the $I=0$ $D\bar{D}^{*}$ scattering is considered. In this sense, the energies $E_{n}$ (the bottom row in Tables SI1-SI4) from both both $c\bar{c}$ and $D\bar{D}^{*}$ operators are taken as good estimates of the eigenvalues of the lattice Hamiltonian relevant to the $I^{G}J^{PC}=0^{+}1^{++}$ charmonium-like system.
@@
Table SI1: Energy levels $E_{n}$ from different subsets of operators for lattice ensemble $m_{\pi}=250~{}\mathrm{MeV}$. In each row of $E_{n}$, the energy values (converted to the physical unit GeV through the fixed $a_{t}^{-1}$ in Table SI5) are obtained from the operators in the non-empty columns. Therefore, the first four rows of $E_{n}$ show the unmixed energies. The mixing effects are very weak and can almost be neglected for $D\bar{D}^{*}(q=0)-D\bar{D}^{*}(q=1)$ (row no.5), $J/\psi\omega-D\bar{D}^{*}$ (row no.6) and $J/\psi\omega-c\bar{c}$ (row no.7) mixings. The strong mixing takes place between $c\bar{c}$ states and $D\bar{D}^{*}$ states, which is manifested by the substantial changes of energies $E_{2}$ and $E_{3}$ after the mixing (row no.8 and row no.9).
@@
Table SI2: Similar to Tab. SI1 but for $m_{\pi}=307~{}\mathrm{MeV}$.
@@
Table SI3: Similar to Tab. SI1 but for $m_{\pi}=362~{}\mathrm{MeV}$.
@@
Table SI4: Similar to Tab. SI1 but for $m_{\pi}=417~{}\mathrm{MeV}$.
@@
Figure SI3: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at various $m_{\pi}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $\{Z_{\alpha}^{(n)},\alpha=1,2,\ldots,6\}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
To be specific, we determine the finite volume energies based on the correlation matrix of $c\bar{c}$ and $D\bar{D}^{*}$ operators, which compose an operator set $\tilde{\mathcal{S}}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,6\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1}\}$. By solving the GEVP to the corresponding correlation matrix $C_{\alpha\beta}(t)$ similarly to Eq. (SI9), we obtain six eigenvectors $\{v_{\alpha}^{(n)},n=1,2,\ldots,6\}$ and six optimized correlation functions $\{C^{(n)}(t)=v_{\alpha}^{(n)}v_{\beta}^{(n)}C_{\alpha\beta}(t),n=1,2,\ldots,6\}$. It is expected that $C^{(n)}(t)$ is contributed most from the $n$-th eigenstate with the eigen-energy $E_{n}$ of the lattice Hamiltonian. The physical status of these states can be explored qualitatively by the coupling factor $Z_{\alpha}^{(n)}$ of $\mathcal{O}_{\alpha}$ to the $n$-th state, which is extracted through the expression

### v2

@@ III Summary
We calculate the $I^{G}J^{PC}=0^{+}1^{++}$ channel $D\bar{D}^{*}$ scattering at four pion masses ranging from 250 MeV to 417 MeV in $N_{f}=2$ lattice QCD. We use the distillation method to calculate the related correlation matrix of $c\bar{c}$ operators and $D\bar{D}^{*}$ operators and obtain three energy $E_{2,3,4}$ levels around the $D\bar{D}^{*}$ threshold. At $m_{\pi}\approx 417$ MeV where there are no severe OPE lhc issues, a normal ERE analysis to the two energy levels $E_{2,3}$ results in a shallow bound state, which resides below the $D\bar{D}^{*}$ by $1.3_{-1.0}^{+0.8}$ MeV and has a compositeness $X\approx 1$ indicating a predominant $D\bar{D}^{*}$ component. The phase shifts $p\cot\delta_{0}(p)$ at $E_{3,4}$ indicate the possible existence of a resonance below $4.0$ GeV.
We also perform a joint analysis of the three energy levels $E_{2,3,4}$ through the $K$-matrix parameterization of the scattering amplitude. After solving the pole equation in the physical Riemann sheet, a genuine bound state is obtained, whose properties are similar to those of the bound state derived through the ERE analysis. This bound state can correspond to $X(3872)$. In addition, we observe a resonance pole in the second Riemann sheet with a mass slightly higher than 4.0 GeV and a width of 40-60 MeV. This resonance decays into $D\bar{D}^{*}$ by almost 100%. It is interesting to note that the properties of this resonance are in fairly good agreement with the newly observed $\chi_{c1}(4010)$ by LHCb Aaij et al. (2024). Anyway, we remark that this result should be checked by more sophisticated lattice QCD calculations in the future with more finite volume energy levels from more lattice volumes and more kinetic frames Prelovsek et al. (2021); Wilson et al. (2024a, b).
Acknowledgements.
@@
We are grateful to Profs. Qiang Zhao, Feng-Kun Guo, Jia-Jun Wu, Zhihui Guo and Xu Feng for the valuable discussions. This work is supported by the National Natural Science Foundation of China (NNSFC) under Grants No. 11935017, No. 12293060, No. 12293061, No. 12293065, No. 12293063, No. 12075253, No. 12070131001 (CRC 110 by DFG and NNSFC), No. 12175063, No. 12205311, the National Key Research and Development Program of China (No. 2020YFA0406400) and the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030302). The Chroma software system Edwards and Joo (2005) and QUDA library Clark et al. (2010); Babich et al. (2011) are acknowledged. The computations were performed on the HPC clusters at the Institute of High Energy Physics (Beijing), China Spallation Neutron Source (Dongguan), and the ORISE computing environment.
Appendix A Conventions and Operators
These appendix provide further information on our study on the $D\bar{D}^{*}(I=0)$ scattering process. The contents include the details of the operators’ construction, the calculation of correlation matrix, and the derivation of finite volume energy levels, etc.
@@ Appendix A Conventions and Operators
$\begin{cases}|u\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |d\rangle=\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}\begin{cases}|\bar{d}\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |\bar{u}\rangle=-\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}$ (17)
@@
$\begin{cases}|D^{(*)+}\rangle=|c\bar{d}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)0}\rangle=|c\bar{u}\rangle=-|\frac{1}{2},-\frac{1}{2}\rangle,\\ |\bar{D}^{(*)0}\rangle=|u\bar{c}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)-}\rangle=|d\bar{c}\rangle=|\frac{1}{2},-\frac{1}{2}\rangle.\\ \end{cases}$ (18)
@@
$|D\bar{D}^{*}\rangle_{I=0}^{Q=0}$ $=$ $\frac{1}{2}\left(|D^{+}\bar{D}^{*-}\rangle+|D^{0}\bar{D}^{*0}\rangle\right.$ (19)
−$\left.-|\bar{D}^{0}D^{*0}\rangle-|D^{-}D^{*+}\rangle\right).$
Since the charge conjugation transformation ($\mathcal{C})$ of $D$ and $D^{*}$ are conventionally defined as $\mathcal{C}|D\rangle=+|\bar{D}\rangle$, $\mathcal{C}|D^{*}\rangle=-|\bar{D}^{*}\rangle$, one can easily check the $\mathcal{C}$ transformation $\mathcal{C}|D\bar{D}^{*}\rangle_{I=0}^{Q=0}=+|D\bar{D}^{*}\rangle_{I=0}^{Q=0}$. The $\mathcal{P}$-parity $P=+$ requires the $D\bar{D}^{*}$ relative angular momentum $l=0$. Therefore, the $S$-wave $D\bar{D}^{*}$ state with the combination in Eq. (19) has the desired quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. The combination in Eq. (19) gives also the flavor structure of the interpolation field operator of $D\bar{D}^{*}$ systems.
A $S$-wave two-particle operator $\mathcal{O}_{AB}(q)$ is easily constructed by summing over all the relative momenta $\mathbf{p}\equiv 2\pi\mathbf{q}/L$ with the same $q=|\mathbf{q}|$, namely,
$\mathcal{O}_{AB}^{q}=\frac{1}{N_{q}}\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q}).$ (20)
where $R$ runs over all the elements of the octahedral group $O$ and $N_{q}$ is the degeneracy of $q$. Given the single-particle operators of $D$ and $\bar{D}^{*}$, $(\mathcal{O}_{D},\mathcal{O}_{D^{*}})\equiv(\bar{u}(\bar{d})\Gamma_{D}c,\bar{u}(\bar{d})\Gamma_{D^{*}}c)$, we combine the relations in Eq. (19) and (20) to build up the $D\bar{D}^{*}$ operators $\mathcal{O}_{D\bar{D}^{*}}^{q}$ with the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. In practice, we use two combinations of $\Gamma_{D}$ and $\Gamma_{D^{*}}$, namely, $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$ and $(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$. The operators $\mathcal{O}_{D\bar{D}^{*}}^{q=0,\gamma_{5}(\gamma_{4}\gamma_{5})}$ with the two $(\Gamma_{D},\Gamma_{D^{*}})$ combinations (indicated by the superscripts $\gamma_{5}$ and $\gamma_{4}\gamma_{5}$) are also named by $\mathcal{O}_{4}$ and $\mathcal{O}_{5}$, respectively. The $\mathcal{O}_{D\bar{D}^{*}}^{q=1}$ operator uses the first $(\Gamma_{D},\Gamma_{D^{*}})$ combination and is also named as $\mathcal{O}_{6}$.
@@
Figure A1: Schematic quark contraction diagrams for the calculation of the correlation matrix $C_{\alpha\beta}(t)$. For convenience, the operator set is sorted as $\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\}$. All the contributions are considered except for charm quark annihilation diagrams which are expected to be suppressed by the OZI rule.
@@
Figure A2: Effective mass using different operators subset and solving GEVP, ranging from different $m_{\pi}$. In each column of the panels, the headers indicate the operators involved in GEVP. The colored lines show the effective mass behavior corresponding to their dominant operators, as indicated by the same color. Their fitting values are given in Table.[A1-A4].
@@
Table A1: Energy levels $E_{n}$ from different subsets of operators for lattice ensemble $m_{\pi}=250~{}\mathrm{MeV}$. In each row of $E_{n}$, the energy values (converted to the physical unit GeV through the fixed $a_{t}^{-1}$ in Table C1) are obtained from the operators in the non-empty columns. Therefore, the first four rows of $E_{n}$ show the unmixed energies. The mixing effects are very weak and can almost be neglected for $D\bar{D}^{*}(q=0)-D\bar{D}^{*}(q=1)$ (row no.5), $J/\psi\omega-D\bar{D}^{*}$ (row no.6) and $J/\psi\omega-c\bar{c}$ (row no.7) mixings. The strong mixing takes place between $c\bar{c}$ states and $D\bar{D}^{*}$ states, which is manifested by the substantial changes of energies $E_{2}$ and $E_{3}$ after the mixing (row no.8 and row no.9).
@@
Table A2: Similar to Tab. A1 but for $m_{\pi}=307~{}\mathrm{MeV}$.
@@
Table A3: Similar to Tab. A1 but for $m_{\pi}=362~{}\mathrm{MeV}$.
@@
Table A4: Similar to Tab. A1 but for $m_{\pi}=417~{}\mathrm{MeV}$.
@@
Appendix B Determination of finite volume energies
Based on the operator set introduced above,
$\mathcal{S}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\},$ (23)
we calculate the corresponding correlation matrix